1. Mullineaux Co. issued 11-year bonds one year ago at a coupon rate of 8.6 percent. The bonds make semiannual payments. If the YTM on these bonds is 7.5 percent, what is the current bond price?
Answer #1


Using financial calculator BA II Plus - Input details:

#

I/Y = R = Rate or yield / frequency of coupon in a year = 7.5/12 =

                    3.750000

PMT = Coupon rate x FV / frequency = 8.6% x -1000 / 2 =

-$43.00

N = Number of years remaining x frequency = Issue 1 year ago = 10 x 2 =

20.00

FV = Future Value =

-$1,000.00

CPT > PV = Present value of bond = Price of Bond = Current bond price =

$1,076.429

Formula for bond value = |PMT| x ((1-((1+R%)^-N)) / R%) + (|FV|/(1+R%)^N) =

$1,076.429